# Memory preservation and cooperative shielding in complex quantum networks

### v2

@@ 4 Discussion and conclusions
Acknowledgements
We would like to thank Concetta Morrone and Marco Cicchini for decisive discussions on the neuroscience application of the analytical findings of this work, specifically concerning the connection between the number of frequencies and the number of excitations, and of the energy bandwith scaling with the Weber’s law. J.Y.M. and M.L.C. were supported by the European Social Fund REACT EU through the Italian national program PON 2014-2020, DM MUR 1062/2021. M.L.C. acknowledges support from the National Centre on HPC, Big Data and Quantum Computing—SPOKE 10 (Quantum Computing) and received funding from the European Union Next-GenerationEU—National Recovery and Resilience Plan (NRRP)—MISSION 4 COMPONENT 2, INVESTMENT N. 1.4—CUP N. I53C22000690001. This research has received funding from the European Union’s Digital Europe Programme DIGIQ under Grant Agreement No. 101084035. M.L.C. also acknowledges support from the project PRA2022202398 “IMAGINATION.” M.L.C, F.B. and G.L.C. are grateful to the KITP for the valuable hospitality during the conception of part of this work. This research was supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP).
@@ Appendix A Degeneracy formula for the XX, all-to-all connected energy levels
In Section 2 we introduced the recursive formula $f(l,L)$ for computing the degeneracy of a certain value of the $\hat{\mathcal{M}}^{2}$ quantum number $l$ for a given system size $L$; here we prove the result. As the Hamiltonian in Eq. 2 is written in terms of global spin operators, the Hilbert space of possible spin configurations (Pauli strings) can be decomposed into irreducible representations of spin $l$, each one with dimension $2l+1$: this means that the total Hilbert space is a direct sum of such representations.
@@
Appendix B Inverse Participation Ratio
@@
Appendix C Characterizing complex graphs: average distance and clustering
@@
The clustering coefficient is instead a measure of the nodes’ tendency to congregate. While there are many possible definitions, we refer to the one given by Watts and Strogatz in [78] since it is defined for the class of graphs we consider in this work, i.e. small-world networks. Specifically, it refers to the global clustering coefficient computed as average of the local ones. Given a node $q$ with $k_{q}$ neighbors (in other words it represents the degree of the site), we define the local clustering coefficient $C_{q}$ as the proportion $t_{q}$ of the number of triangles, i.e. the closed loops consisting of 3 nodes, containing $q$, counted with respect to the maximum number of allowed connections $k_{q}(k_{q}-1)/2$ (if $k_{q}$ is greater than 1): the latter is also referred in the literature as the overall amount of triplets, or subgraphs composed of three nodes connected with at least two edges. Using the properties of the graph’s adjacency matrix $A$ ($A_{ij}=1$ if vertices $i$ and $j$ are connected, 0 otherwise), specifically the one that says that its powers indicate the number of walks between two sites, one can compute $t_{q}$ as the matrix element $(A^{3})_{qq}$ divided by 2 (graphs are undirected, so this allows to count the unique triangles). In summary, the clustering coefficient $\mathcal{C}$ comes as follows:
$t_{q}=\frac{(A^{3})_{qq}}{2}\Longrightarrow C_{q}=\begin{cases}0\quad\text{if $k_{q}=0,1$}\\ t_{q}/[k_{q}(k_{q}-1)/2]\quad\text{if $k_{q}>1$}\end{cases}\Longrightarrow\mathcal{C}=\frac{1}{L}\sum_{q}C_{q}\,.$

### v3

@@ IV Discussion and conclusions
Acknowledgements.
We would like to thank Concetta Morrone and Marco Cicchini for decisive discussions on the neuroscience application of the analytical findings of this work, specifically concerning the connection between the number of frequencies and the number of excitations, and of the energy bandwith scaling with the Weber’s law. S.A. gratefully thanks École polytechnique and Karyn Le Hur, together with the Quantum Sensing Group, Samuel Lellouch and Michael Holynski, for allowing the fulfillment of the present research. J.Y.M. and M.L.C. were supported by the European Social Fund REACT EU through the Italian national program PON 2014-2020, DM MUR 1062/2021. M.L.C. acknowledges support from the National Centre on HPC, Big Data and Quantum Computing—SPOKE 10 (Quantum Computing) and received funding from the European Union Next-GenerationEU—National Recovery and Resilience Plan (NRRP)—MISSION 4 COMPONENT 2, INVESTMENT N. 1.4—CUP N. I53C22000690001. This research has received funding from the European Union’s Digital Europe Programme DIGIQ under Grant Agreement No. 101084035. M.L.C. also acknowledges support from the project PRA2022202398 “IMAGINATION.” M.L.C, F.B. and G.L.C. are grateful to the KITP for the valuable hospitality during the conception of part of this work. This research was supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP).
@@ Appendix A Degeneracy formula for the XX, all-to-all connected energy levels
In Section II we introduced the recursive formula $f(l,L)$ for computing the degeneracy of a certain value of the $\hat{\mathcal{M}}^{2}$ quantum number $l$ for a given system size $L$; here we prove the result. As the Hamiltonian in Eq.(2) is written in terms of global spin operators, the Hilbert space of possible spin configurations (Pauli strings) can be decomposed into irreducible representations of spin $l$, each one with dimension $2l+1$: this means that the total Hilbert space is a direct sum of such representations.
@@
Appendix B Numerical methods
Throughout this work we simulated the systems here reported using Exact Diagonalization (ED) techniques: while the errors contained in these methods are limited to approximations made in the algorithms employed (namely Krylov subspace methods to implement the unitary time-evolution), their main disadvantage relies on the exponential scaling of computational resources with the chain size. There are numerical variational strategies based on Matrix Product States (MPS) and Density Matrix Renormalization Group (DMRG) [schollwoeck2011mps] that have met great success in simulating strongly-correlated quantum systems. However, we employed ED for simulating the quantum dynamics of both all-to-all and regular networks because of the exponential complexity in the number of time steps associated with MPS/DMRG methods.
We are able to circumvent the problem of system size exponential complexity of the ED method thanks to the Hilbert space restriction to symmetry sectors. Indeed, through the computational basis introduced in Section II, the Hamiltonian (1) can be block-diagonalized, each block being defined by the set of eigenstates with a given total $z$-magnetization of the state. This symmetry implies that if we initially consider a state with a well-defined magnetization, our computation of the dynamics simplifies enormously, as it reduces the size of the operators’ matrices that we need to compute: for example, by introducing initial states possessing an initial magnetization of m=(L/2)+1m=(-L/2)+1 (meaning all spins down except for one) the subspace of relevance will be of size $L$ (binomial coefficient “$L$ choose 1”) instead of $2^{L}$.
Appendix C Inverse Participation Ratio
@@
Appendix D Characterizing complex graphs: average distance and clustering
@@
The clustering coefficient is instead a measure of the nodes’ tendency to congregate. While there are many possible definitions, we refer to the one given by Watts and Strogatz in [watts1998smallworld] since it is defined for the class of graphs we consider in this work, i.e. small-world networks. Specifically, it refers to the global clustering coefficient computed as average of the local ones. Given a node $q$ with a number of neighbors given by the degree degq, we define the local clustering coefficient $C_{q}$ as the proportion $t_{q}$ of the number of triangles, i.e. the closed loops consisting of 3 nodes, containing $q$, counted with respect to the maximum number of allowed connections between node $q$ and its neighbors, $\text{deg}_{q}(\text{deg}_{q}-1)/2$ (if degq is greater than 1): the latter is also referred in the literature as the overall amount of triplets, or subgraphs composed of three nodes connected with at least two edges. Using the properties of the graph’s adjacency matrix $A$ ($A_{ij}=1$ if vertices $i$ and $j$ are connected, 0 otherwise), specifically the one that says that its powers indicate the number of walks between two sites, one can compute $t_{q}$ as the matrix element $(A^{3})_{qq}$ divided by 2 (graphs are undirected, so this allows to count the unique triangles). In summary, the clustering coefficient $\mathcal{C}$ comes as follows:
$t_{q}$ $=\frac{(A^{3})_{qq}}{2}$
[…]
$\mathcal{C}$ $=\frac{1}{L}\sum_{q}C_{q}\,.$